# Distributed memory Tensor-train crossT. Shi, D. Hayes, and J. Qiu

### v1

@@ 4 Process grid development for subtensor TT cross
For simplicity, for a $d$-dimensional tensor $\mathcal{X}$, we assume it has uniform mode size $n$, subtensor partitioning count per dimension $C$, and process count per dimension $P$. Note that unlike in Section 3 that we assume $P=C$, in this section we no longer have this requirement. In general, one may consider process grids partition the distributed memory machines, while subtensor grids partition the data into optimal size for handling and algorithm performance. In practice, subtensor partitioning is often a refinement of process partitioning. In addition, we also assume the TT core size is $r$ for all dimensions, so the number of pivots we select for each dimension is $r$ as well. Suppose $n=Cm$, then $m$ is the mode size per dimension of the subtensor. If there are not enough processes to fill a $d$-dimensional process grid, it’s easy to see that one can apply any $z$-dimensional grid as long as $z<d$. In this case, mode sizes of certain dimensions of the subtensors are multiples of $m$. However, lower dimensional process grids can lead to load imbalance issues when $C$ is not a multiple of $P$. In this section, we assume $C=Pw$, so that each process handles $w^{d}$ subtensors.
As remarked in Section 3.2Algorithm 4 has two sets of process grids for Algorithm 2 and Algorithm 3: (a) $d$-dimensional grid for pivot selection, and (b) $d-1$ lower dimensional grids for TT core construction. Notice that processes in Grid (b) also store tensor elements, and are chosen directly from Grid (a).
@@
  • Grid (a): $d$-dimensional grid for pivots. In TT pivot selection, Algorithm 2 is used $d-1$ times, one for each tensor unfolding. In this step, we need to study both the Allgather for pivots, and communications with row and column neighbors. Once we find local pivots on all subtensors, we can compare among the $w^{d}$ subtensors on the same process to get $P^{d}$ potential new pivots for each dimension, labeled by $(i_{\leq j}^{(p)},i_{>j}^{(p)})$. Here, $11\leq j\leq d-$ 1 is used to denote dimensions, and $1\leq p\leq P^{d}$ is used to denote a specific process. Next, Allgather allows all processes to know all pivots $(i_{\leq j}^{(p)},i_{>j}^{(p)})$, together with the pivot selection metric. This procedure requires each process to send $3(P^{d}-1)$ elements for all dimensions, and receive a throughput of $3(P^{d}-1)$ entries. In this way, the best global pivot for a dimension $k$ can be found on each process.
@@
    elements, where the two terms are the vectors received from column and row neighbors respectively. Comparatively, the process with the pivot needs to send the most information to its row and column neighbors, of an approximate entry count
@@
  • Connecting Grid (a) and Grid (b). The gap between Grid (a) and Grid (b) is bridged by lines 5-10 in Algorithm 4. Lines 5-9 communicate about the pivots with grids for adjacent dimensions, and line 10 get necessary tensor entries for construction. Communications across the lower dimensional grids in lines 5-9 is straightforward, as they only occur for the pivot index sets on the grid roots. For dimension $j$, this means sending $r$ values to dimension $j+1$, and receiving $r$ values from dimension $j-1$. The analysis of line 10 is a bit more complicated. In the extreme case that one process does not contain any of the pivots, it needs to obtain tensor elements associated with the subtensors $\mathcal{X}(\mathcal{I}_{\leq j},\mathcal{J}_{>j})$ and $\mathcal{X}(\mathcal{K}_{1},\cdots,\mathcal{K}_{j},\mathcal{J}_{>j})$ for $11\leq j\leq d-$ 1. Therefore, a pessimistic bound of the total number of elements received is
@@
  • Grid (b): Lower-dimensional grids for TT construction. For a $d$-way tensor, Grid (b) contains $d-1$ lower-dimensional grids for the row indices of tensor unfoldings, i.e. $j$-dimensional grid for the $j$th unfolding for $11\leq j\leq d-$ 1, as the row indices of this unfolding correspond to $j$ dimensions. The communications only happen in line 11 of Algorithm 4, within the $j$-dimensional grid. Since one process can only appear once in one of the lower dimensional grids, the bound Eq. 12 can be applied to the analysis as well. As the final TT core is simply a reshape of the computed portion of $T_{\leq j}$, this part of the algorithm does not involve a lot of hanging procedures and communications, and is thus very fast in practice.
@@ 5 Numerical Examples
In this section, we show the performance of our parallel TT cross algorithm [^1] The code for implementation can be found at github.com/dhayes95/Cross. on the Hilbert tensors (see Section 5.2) and discretized Maxwellian functions as equilibrium distribution functions in kinetic theory [cercignani1969mathematical] (see Section 5.3). In all numerical tests, results are computed on the DARWIN (Delaware Advanced Research Workforce and Innovation Network) system. The standard partition of this system consists of 32-core AMD EPYC 7002 2.80 GHz series processors. In addition, for all examples, any mention of error measurement corresponds to the relative error on a uniformly generated set of $R$ random indices $i$
$||\mathcal{X}-\tilde{\mathcal{X}}||_{F,R}^{2}:=\frac{\sum^{R}{(\mathcal{X}(i)-\tilde{\mathcal{X}}(i)})^{2}}{\sum^{R}{\mathcal{X}(i)^{2}}}.$
This error approximation allows us to visualize the accuracy of our solver without going through all the elements of a large dataset.
@@ 5.1 Testing setup
For all results presented, the algorithm is performed for 10 runs, and the minimum times and errors are taken for plots. For the error plots, all core ranks are incrementally increased until they reach the specified core rank used; e.g., in the $4$-d Maxwellian test in Section 5.3 with prescribed rank $(1,10,5,20,1)$, the core ranks used for error tests are
@@ 5.2 Hilbert tensor
where we assume a uniform mode size $n$. It is known that these Hilbert tensors can be accurately approximated by a numerically low TT rank tensor, and the TT ranks $r$ can be estimated a priori given $d$ and $n$ [shi2021compressibility]. We shall use this rank approximation as the number of pivots we select in our tests.
First, we test the case of $d=3$ with mode size $n=2000$ and core ranks $(1,25,25,1)$, and refer to Table 1 to see partitioning set up for certain MPI ranks in scaling tests. In Fig. 7, we can see the plots for strong and weak scaling results. This test shows good strong scaling results as the pivot selection demonstrates almost optimal results. For the core construction we see for 1-8 MPI ranks a larger slope. The cause for this phenomena has not currently been investigated, but will be our future target. Following this, we see the scaling from 8-64 MPI ranks demonstrate almost optimal linear scaling. For the weak scaling test, we increase both the mode size and the number of processes linearly, while maintaining a subtensor of size $500\times 500\times 500$ on each process. In Fig. 7 we can see in the middle plot that as the number of processes increases we do not suffer from large growth in computational time for pivot selection nor core construction. As with the strong scaling results, the cause behind small intermittent decrease in time is not currently investigated. We also include results of sampled errors in Fig. 7 (Right). Here the horizontal axis corresponds to the core ranks $(1,r,r,1)$ and the vertical axis are the values $||\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$. This plot shows that we are able to achieve a close approximation of the true tensor value in the distributed memory setting.
Figure 7: Hilbert tensor with $d=3$ and $n=2000$. Left: strong scaling. Middle: weak scaling. Right: accuracy approximation with $||\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$
@@
Figure 8: Hilbert tensor with $d=6$ and $n=300$. Top left: strong scaling. Top right: weak scaling. Bottom left: accuracy approximation with $||\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$. Bottom right: percentage of full tensor size accessed to perform construction.
@@ 5.3 Tensors from Maxwellian distribution functions
In this section, we show two example tensors, with dimension 4 and dimension 6, constructed from distribution functions arise from kinetic theory of gas dynamics. In both cases, the tensors are constructed element-wise from function values on a discretized grid in both displacement and velocity spaces. For the $4$-dimensional (2d2v) tensor, the function used to compute entries of $\mathcal{X}$ is given by
@@
In our tests, the underlying domains are taken to be $[-1/2,1/2]$ for $x,y$ and $[-3,3]$ for $v_{x},v_{y}$. Then our entries of $\mathcal{X}$ are defined by
@@
For the first test, we work with 2d2v with size $(2n,n,2n,n)$ where $n=1000$ and core ranks $(1,10,5,20,1)$. As with the Hilbert tensor, the MPI partitioning can be found in Table 2Fig. 9 shows the plots for both strong and weak scaling for the 2d2v test case. We observe good strong scaling results for all partitions. The weak scaling results display similar results to the Hilbert tensor tests. We also include an error plot of $||\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$, which verifies that we can get a close approximation even with variation in the core ranks as seen in Fig. 9.
Figure 9: Maxwellian tensor with $d=4$. Left: strong scaling. Middle: weak scaling. Right: Accuracy approximation with $||\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$
@@
Figure 10: Maxwellian tensor with $d=6$. Top left: strong scaling. Top right: weak scaling. Bottom left: Accuracy approximation with $||\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$. Bottom right: Percentage of full tensor size accessed to perform construction.

### v2

@@ 4 Process grid development for subtensor TT cross
For simplicity, for a $d$-dimensional tensor $\mathcal{X}$, we assume it has uniform mode size $n$, subtensor partitioning count per dimension $C$, and process count per dimension $P$. Note that unlike in Section 3 that we assume $P=C$, in this section we no longer have this requirement. In general, one may consider process grids partition the distributed memory machines, while subtensor grids partition the data into optimal size for handling and algorithm performance. In practice, subtensor partitioning is often a refinement of process partitioning. In addition, we also assume the TT core size is $r$ for all dimensions, so the number of pivots we select for each dimension is $r$ as well. Suppose $n=Cm$, then $m$ is the mode size per dimension of the subtensor. If there are not enough processes to fill a $d$-dimensional process grid, it’s easy to see that one can apply any $z$-dimensional grid as long as $z<d$. In this case, mode sizes of certain dimensions of the subtensors are multiples of $m$. However, lower dimensional process grids can lead to load imbalance issues when $C$ is not a multiple of $P$, since the subtensors cannot have uniform mode size. As a result, some processes compute with more subtensors, or subtensors with larger size. In this section, we assume $C=Pw$, so that each process handles $w^{d}$ subtensors, and the number of elements per dimension held by one process is $wm$.
As remarked in Section 3.2Algorithm 5 has two sets of process grids for Algorithm 3 and Algorithm 4: (a) $d$-dimensional grid for pivot selection, where each process in the grid stores subtensors of the target tensor, and (b) d1d-1 lower dimensional grids for TT core construction, where each process is only responsible to generate part of one TT core, and thus needs to store the tensor elements corresponding to the used pivots in Algorithm 4. As the process is already part of grid (a) and stores some tensor entries, it needs to receive the necessary data from other processes in grid (a).
@@
  • Grid (a): $d$-dimensional grid for pivots. In TT pivot selection, Algorithm 3 is used d1d-1 times, one for each tensor unfolding. In this step, we need to study both the Allgather for pivots, and communications with row and column neighbors. Once we find local pivots on all subtensors, we can compare among the $w^{d}$ subtensors on the same process to get $P^{d}$ potential new pivots for each dimension, labeled by $(i_{\leq j}^{(p)},i_{>j}^{(p)})$. Here, $1\leq j\leq d-1$ is used to denote dimensions, and $1\leq p\leq P^{d}$ is used to denote a specific process. Next, Allgather allows all processes to know all pivots $(i_{\leq j}^{(p)},i_{>j}^{(p)})$, together with the pivot selection metric. This procedure requires each process to send $3(P^{d}-1)$ elements for all dimensions, and receive a throughput of $3(P^{d}-1)$ entries. In this counting, 3 consists of the two indices of the local pivot and the pivot value, and $P^{d}-1$ is the total number of other processes to communicate to get the global pivot. In this way, the best global pivot for a dimension $k$ can be found on each process.
@@
    elements, where the two terms are the vectors received from column and row neighbors respectively (see Fig. 3). Comparatively, the process with the pivot needs to send the most information to its $P^{j}-1$ row and $P^{d-j}-1$ column neighbors, of an approximate entry count
@@
  • Connecting Grid (a) and Grid (b). The gap between Grid (a) and Grid (b) is bridged by lines 5-10 in Algorithm 5. Lines 5-9 communicate about the pivots with grids for adjacent dimensions, and line 10 get necessary tensor entries for construction. Communications across the lower dimensional grids in lines 5-9 is straightforward, as they only occur for the pivot index sets on the grid roots. For dimension $j$, this means sending $r$ values to dimension $j+1$, and receiving $r$ values from dimension j1j-1. The analysis of line 10 is a bit more complicated. In the extreme case that one process does not contain any of the pivots, it needs to obtain tensor elements associated with the subtensors $\mathcal{X}(\mathcal{I}_{\leq j},\mathcal{J}_{>j})$ and $\mathcal{X}(\mathcal{K}_{1},\cdots,\mathcal{K}_{j},\mathcal{J}_{>j})$ for $1\leq j\leq d-1$ , which contains $r^{2}$ and $r(wm)^{j}$ elements respectively. Therefore, a pessimistic bound of the total number of elements received is
@@
  • Grid (b): Lower-dimensional grids for TT construction. For a $d$-way tensor, Grid (b) contains d1d-1 lower-dimensional grids for the row indices of tensor unfoldings, i.e. $j$-dimensional grid for the $j$th unfolding for $1\leq j\leq d-1$, as the row indices of this unfolding correspond to $j$ dimensions. The communications only happen in line 11 of Algorithm 5, within the $j$-dimensional grid. Since one process can only appear once in one of the lower dimensional grids, the bound Eq. 13 can be applied to the analysis as well. As the final TT core is simply a reshape of the computed portion of $T_{\leq j}$, this part of the algorithm does not involve a lot of hanging procedures and communications, and is thus very fast in practice.
@@ 5 Numerical Examples
In this section, we show the performance of our parallel TT cross algorithm [^1] The code for implementation can be found at github.com/dhayes95/Cross. on the Hilbert tensors (see Section 5.2) and discretized Maxwellian functions as equilibrium distribution functions in kinetic theory [10, 50] (see Section 5.3). For all tests, the tensors used are represented element-wise using function evaluations. This allows for low memory requirement as the construction used does not require full tensor data, so only entries that are required for construction are computed on the fly.
In all numerical tests, results are computed on the DARWIN (Delaware Advanced Research Workforce and Innovation Network) system. The standard partition of this system consists of two 32-core AMD EPYC 7002 2.80 GHz series processors allowing for 64 total cores per compute node. At the current time, DARWIN does not support node-to-node communication using MPI4PY, so a distributed memory setting is simulated by giving each CPU its own memory. In addition, for all examples, any mention of error measurement corresponds to the relative error on a uniformly generated set of $R$ random indices $i$
$\mathcal{X}-\tilde{\mathcal{X}}||_{F,R}^{2}:=\frac{\sum^{R}{(\mathcal{X}(i)-\tilde{\mathcal{X}}(i)})^{2}}{\sum^{R}{\mathcal{X}(i)^{2}}}.$
This error approximation allows us to visualize the accuracy of our solver without having to go through all the elements of a large dataset.
@@ 5.1 Testing setup
In all test cases, the core ranks used are selected experimentally with the cardinality of the pivot sets in ACA. While the Hilbert tensor has theoretical justification for the core ranks [44], core ranks of the Maxwellian tensors are theoretically unknown and require the experimental selection. For all results presented, the algorithm is performed for 10 runs, and the minimum times and errors are taken for plots. For the error plots, all core ranks are incrementally increased until they reach the specified core rank used; e.g., in the $4$-d Maxwellian test in Section 5.3 with prescribed rank $(1,10,5,20,1)$, the core ranks used for error tests are
@@ 5.2 Hilbert tensor
where we assume a uniform mode size $n$. It is known that these Hilbert tensors can be accurately approximated by a numerically low TT rank tensor, and the TT ranks $r$ can be estimated a priori given $d$ and $n$ [44]. We shall use this rank approximation as the number of pivots we select in our tests.
First, we test the case of $d=3$ with mode size $n=2000$ and core ranks $(1,25,25,1)$, and refer to Table 2 to see partitioning set up for certain MPI ranks in scaling tests. In Fig. 8, we can see the plots for strong and weak scaling results. This test shows good strong scaling results as the pivot selection demonstrates almost optimal results, where the optimal line corresponds to a line with slope 1, i.e. time is cut in half when the number of MPI ranks is doubled. For the core construction we see for 1-8 MPI ranks a larger slope. The cause for this phenomena has not currently been investigated, but will be our future target. Following this, we see the scaling from 8-64 MPI ranks demonstrate almost optimal linear scaling. For the weak scaling test, we increase both the mode size and the number of processes linearly, while maintaining a subtensor of size $500\times 500\times 500$ on each process. In Fig. 8 we can see in the middle plot that as the number of processes increases we do not suffer from large growth in computational time for pivot selection nor core construction. As with the strong scaling results, the cause behind small intermittent decrease in time is not currently investigated. We also include results of sampled errors in Fig. 8 (Right). Here the horizontal axis corresponds to the core ranks $(1,r,r,1)$ and the vertical axis are the values $\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$. The core ranks are selected to be $(1,r,r,1)$ as $3D$ Hilbert tensors admit symmetric core ranks from Eq. 19. This plot shows that we are able to achieve a close approximation of the true tensor value in the distributed memory setting.
Figure 8: Hilbert tensor with $d=3$ and $n=2000$. Left: strong scaling. Middle: weak scaling. Right: accuracy approximation with $\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$ using $64$ MPI ranks.
@@
Figure 9: Hilbert tensor with $d=6$ and $n=300$. Top left: strong scaling. Top right: weak scaling. Bottom left: accuracy approximation with $\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$. Bottom right: percentage of full tensor size accessed to perform construction.
@@ 5.3 Tensors from Maxwellian distribution functions
In this section, we show two example tensors, with dimension 4 and dimension 6, constructed from distribution functions arise from kinetic theory of gas dynamics [10, 50]. In both cases, the tensors are constructed element-wise from function values on a discretized grid in both displacement and velocity spaces. For the $4$-dimensional (2d2v) tensor, the function used to compute entries of $\mathcal{X}$ is given by
@@
In our tests, the underlying domains are taken to be [1/2,1/2][-1/2,1/2] for $x,y$ and [3,3][-3,3] for $v_{x},v_{y}$. Then our entries of $\mathcal{X}$ are defined by
@@
For the first test, we work with 2d2v with size $(2n,n,2n,n)$ where $n=1000$ and core ranks $(1,10,5,20,1)$. As with the Hilbert tensor, the MPI partitioning can be found in Table 3Fig. 10 shows the plots for both strong and weak scaling for the 2d2v test case. We observe good strong scaling results for all partitions. The weak scaling results display similar results to the Hilbert tensor tests. We also include an error plot of $\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$, which verifies that we can get a close approximation even with variation in the core ranks as seen in Fig. 10.
Figure 10: Maxwellian tensor with $d=4$. Left: strong scaling. Middle: weak scaling. Right: Accuracy approximation with $\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$ using 64 MPI ranks.
@@
Figure 11: Maxwellian tensor with $d=6$. Top left: strong scaling. Top right: weak scaling. Bottom left: Accuracy approximation with $\mathcal{X}-\tilde{\mathcal{X}}||_{F,10^{6}}$. Bottom right: Percentage of full tensor size accessed to perform construction.
5.4 Comparisons with dimension parallel and torchTT
In the first part of this subsection, we make a comparison of timings for pivot selection of our proposed algorithm with our own implementation of the dimension parallel TT-Cross algorithm in [17]. The main idea of the dimension parallel algorithm is to select pivots across all dimensions simultaneously, and then communicate indices to neighboring dimensions to update superblocks for the next iteration of searching. The timings reported are in the fully parallel regime of both algorithms for all tensors tested in previous sections. Note that the compression of both algorithms is the same as each is prescribed the same core ranks. Furthermore, the dimension parallel algorithm reported here is only in parallel for index selection, and not core construction.
As seen in Table 4, in both Hilbert tensors tested in Section 5.2 as well as the Maxwellian tensors in Section 5.3, the subtensor parallel algorithm using 64 MPI ranks has significantly smaller run time for index selection compared to the dimension parallel algorithm. These results are expected as the mode size of the tensors are quite large, and the subtensor parallel algorithm uses more computing resources.
Alg./Tensor Hilbert $3d$ Hilbert $6d$ Maxwellian $4d$ Maxwellian $6d$
Subtensor $99.99$ $260.72$ $147.34$ $184.16$
Dimension $2482.75$ $950.43$ $1743.47$ $538.54$
Table 4: Timings in seconds for pivot selection of subtensor parallel compared to dimension parallel in a fully parallel regime. The mode sizes for the various tests are: Hilbert 3dd: (2000,2000,2000), Hilbert 6dd: (300,300,300,300,300,300), Maxwellian 4dd: (2000,1000,2000,1000), Maxwellian 6dd: (800,400,800,400,800,400).
We also compare timings of subtensor parallel and dimension parallel on a Hilbert tensor as dimension increases in two more settings: (1) the number of entries is held approximately constant, and (2) the mode size is fixed. For the first case, the corresponding mode sizes $n$ to dimension $d$ are: $n=1400$ for $d=5$, $n=419$ for $d=6$, $n=177$ for $d=7$, $n=93$ for $d=8$, $n=56$ for $d=9$, and $n=37$ for $d=10$, so that the all of the test cases have roughly $5.3e15$ entries. The choice of total size $5.3e13$ is selected as to run the largest test without running into memory request errors. In the second case, we fix $n=100$ for all cases of dimension $d$.
Figure 12: Timings in seconds for index selection using subtensor parallel and dimension parallel. Left: Constant tensor size with varying mode size. Right: Constant mode size with varying dimension.
As shown in both plots of Fig. 12, for the case of lower dimensions we observe a significant reduction in run time for index selection. Both tests also demonstrate that when dimension grows, the subtensor parallel algorithm approaches the run time of the dimension parallel algorithm. This behavior is expected as the dimension parallel algorithm is designed for tensors with large dimensionality and relatively smaller mode size. Therefore, as the fundamental framework of the subtensor and dimension parallel algorithms are distinct, we are unsurprised to see a transition point where the two perform similarly.
In the last part of this section, we compare the subtensor parallel algorithm with the publicly available python package torchTT [30], which contains a TT-Cross function. In their implementation, they do not use a greedy search such as Algorithm 1, but rather a maxvol search [41]. For our implementation of torchTT, we use a pure python execution that runs in serial. In this comparison, we only test on a $3d$ Hilbert tensor, as the $6d$ Hilbert tensor encounters insufficient memory errors when running torchTT. For $4d$ and $6d$ Maxwellian tensors, the observed time to run torchTT on these tensors significantly exceeds the times achieved by the subtensor algorithms. In a small test case with mode size $n=20$, torchTT takes roughly $1200$ seconds to compress a $6d$ Maxwellian. For this reason, we choose to not attempt a comparison for the other tensors considered in this section.
Alg/$n$ Subtensor $T$-ACA torchTT
Time (s) Error Time (s) Error Time (s) Error
$250$ 1.71 1.27e-06 39.02 1.40e-06 215.34 6.35e-07
$500$ 5.06 9.99e-07 168.09 5.45e-07 860.31 2.93e-06
$1000$ 16.07 9.07e-07 799.42 1.33e-06 3560.26 2.27e-06
$2000$ 60.54 8.64e-07 3605.90 3.76e-07 25159.67 3.19e-06
Table 5: Timings in seconds and error for full approximation using subtensor parallel with 64 MPI ranks, $T$-ACA (subtensor parallel with one process), and torchTT for a $3d$ Hilbert tensor with mode size $n$.
As shown in Table 5, the time taken by the subtensor algorithm is a substantial reduction compared to the times taken by torchTT. It is suspected that one contributing factor to the increase in time is the selection procedure. While maxvol is a robust algorithm for index selection, it is more computationally intensive compared to Algorithm 1, and the accuracy comparison in Table 5 does not show much difference between the two schemes. If we assume that torchTT is implemented with perfect parallel scaling, then with the same number of processes used for the subtensor results of Table 5, torchTT would still report times ranging from $3.36$ seconds to $393.12$ seconds. Therefore, even in a perfect scenario, we still observe substantial reduction in computational time.